OpenStudy (anonymous):

What is the equation of the line whose y-intercept is 3 and slope is 1? y = x - 3 y = x + 3 y = 3x + 1

OpenStudy (anonymous):

y = 3x + 1 as y-intercept is 3 and slope is 1

OpenStudy (aravindg):

@CoreyContini the qns u r asking are too easy ...have u honestly tried them?

OpenStudy (anonymous):

I agree with Aravind, but Rohangrr is wrong, the right answer is y = x + 3

OpenStudy (aravindg):

y=x+3 is the right one

OpenStudy (anonymous):

Every line can be written is the form: y = Ax + B A is your slope, (0,B) is your y-axis intercept

OpenStudy (anonymous):



OpenStudy (anonymous):

as the line rises from left to right it clearly shows that the gradient / slope will be positive. c= where the line intercepts the y axis so y=mx +c m= gradient plug in the values y=x+3

OpenStudy (anonymous):